A D.C wheatstone bridge has a 6.0V supply connected between point A and C what will be the potential difference between point B and D when the resistance in the bridge arms are AB 10(ohms), BC(20ohms), CD(60ohms), AD(31ohms)
1
Expert's answer
2019-10-31T07:30:21-0400

Draw the situation:



According to Kirchoff's laws we can write:


"I=I_1+I_2,\\
I_1(10+20)=6,\\
I_2(60+31)=6.\\
I_1=0.2\text{ A},\\
I_2=0.066\text{ A},\\
I=0.266\text{ A}.\\"

Now imagine that we put a voltmeter between B and D. Then, using the same Kirchoff's laws:


"V_{BD}=V_D-V_B=\\
=(0.066\cdot31-0.066\cdot60)-(0.2\cdot10-0.2\cdot20)=\\
=0.086\text{ V}."